# arXiv:2603.02592v1[math.AG] 03 Mar 2026

### v1

@@ 1 Introduction
Ulrich bundles occupy a distinguished position in the geometry of projective embeddings. They may be viewed as extremal objects with respect to cohomological vanishing and linear resolutions. They play an important role in the study of syzygies, Chow forms, and derived categories [2, 6]. A central problem is to determine which polarized varieties admit Ulrich bundles and to understand their geometric structure when they exist.
The existence question has been resolved in several important cases, including hypersurfaces, complete intersections, and certain classes of surfaces. In particular, Beauville proved the existence of Ulrich bundles on abelian surfaces [1].In higher dimensions, however, systematic constructions remain comparatively sparse, even for classical varieties such as Jacobians.
Beauville’s construction exploits special features of abelian surfaces and does not extend formally to higher-dimensional Jacobians. The present work provides a uniform construction in all dimensions, governed by slope positivity on the curve and Fourier–Mukai duality. In particular, the resulting Ulrich bundles arise functorially from stable bundles on $C$. Our construction produces positive-dimensional families in moduli space, linking the geometry of vector bundles on $C$ to Ulrich theory on $J(C)$. To the best of our knowledge, this provides the first general construction of Ulrich bundles on Jacobians in arbitrary dimension, together with a geometric description of the induced families in moduli.
Let $C$ be a smooth complex projective curve of genus $g>1$, and let
$A=J(C)$
be its Jacobian endowed with its principal polarization $\Theta$. Although the geometry of $A$ is well understood, a uniform construction of Ulrich bundles on the polarized varieties $(A,m\Theta)$ has not previously been carried out in this generality.
The purpose of this paper is to provide such a construction. Our starting point is a semistable vector bundle $V$ on $C$ satisfying the slope condition
$\mu(V)>2g-2.$
Let
$a:C\hookrightarrow A$
be the Abel–Jacobi embedding and let $\mathcal{P}$ denote the normalized Poincaré bundle on $A\times\widehat{A}$. We consider the Fourier–Mukai transform (see subsection 2.3)
$E:=\Phi_{\mathcal{P}}(a_{*}V).$
The key observation is that positivity of slope on the curve translates into strong cohomological control on the Jacobian. More precisely, the inequality $\mu(V)>2g-2$ implies uniform vanishing of $H^{1}(C,V\otimes M)$ for all $M\in Pic^{0}(C)$. Via derived base change, this yields the $\mathrm{WIT}_{0}$ property for $a_{*}V$ and shows that $E$ is locally free with explicitly computable rank. We then prove that the twist $E(\Theta)$ satisfies the stronger $\mathrm{IT}_{0}$ property by reducing cohomology on the Jacobian directly to cohomology on the curve.
The $\mathrm{IT}_{0}$ property implies continuous global generation in the sense of Pareschi–Popa [7]. A kernel construction then produces Ulrich bundles on $(A,m\Theta)$ for all sufficiently positive polarizations.
@@
The rank of the resulting Ulrich bundle is determined by $h^{0}(C,V)$. Specializing to line bundles of degree >2g2>2g-2 yields an upper bound for the Ulrich complexity of $(J(C),m\Theta)$.
Beyond existence, we analyze structural properties of the constructed bundles. Under natural stability assumptions on $V$, we prove that the resulting Ulrich bundles are slope-stable for large polarizations. Using Grothendieck–Riemann–Roch, we compute their Chern classes and show that all bundles arising in the construction have first Chern class proportional to the principal polarization $\Theta$.
The construction is compatible with families and has geometric consequences at the level of moduli space. Fix integers $r\geq 2$ and d>r(2g2)d>r(2g-2), and let $M_{C}$ denote the moduli space of stable vector bundles on $C$ of rank $r$ and degree $d$. The above construction induces a natural morphism from a generically finite cover of $M_{C}$ to the moduli space of stable bundles on $A$ with the corresponding numerical invariants.
This compatibility with families has strong geometric consequences. In particular, the Ulrich locus inside the moduli space of bundles on A is highly nontrivial:
@@
Fix integers $r\geq 2$ and d>r(2g2)d>r(2g-2), and let $M_{C}$ denote the moduli space of stable vector bundles on $C$ of rank $r$ and degree $d$. For $m\geq g+1$, let $M_{A}$ be the moduli space of $H$-stable bundles on $(A,m\Theta)$ with the numerical invariants of the Ulrich bundles constructed in Theorem 5.2 Then there exists a quasi-projective variety $T$, a generically finite surjective morphism
@@
Thus the existence of Ulrich bundles on $(A,m\Theta)$ is governed by slope positivity on the curve and gives rise to a geometric link between vector bundles on $C$ and stable bundles on its Jacobian.
Layout of the paper. Section 2 reviews background on slope stability and Fourier–Mukai transforms. Section 3 establishes the $\mathrm{WIT}_{0}$ property and computes the rank of the transform. Section 4 proves that $E(\Theta)$ satisfies $\mathrm{IT}_{0}$. Section 5 constructs Ulrich bundles and proves Theorem A. Section 6 discusses the numerical bound $m\geq g+1$. Section 7 studies slope stability. Section 8 computes Chern classes. Section 9 specializes to line bundles and derives an upper bound for Ulrich complexity. Section 10 constructs the moduli morphism and proves Theorem B.
@@ 2.2 The Abel–Jacobi embedding
A classical result (see [4, Ch. VII]) states that
@@ 2.7 Semihomogeneous bundles on abelian varieties
  • A semihomogeneous bundle is simple if and only if it is stable with respect to any polarization [15, Prop. 6.16].
  • If $\mathcal{E}$ is semihomogeneous, then its Chern classes are proportional to powers of $c_{1}(\mathcal{E})$ in the Néron–Severi group [15, Thm. 5.8].
@@ 5.1 Positivity and higher cohomology vanishing
For $k\geq 1$, the line bundle $\mathcal{O}_{A}(k\Theta)$ is ample. On an abelian variety, twisting an IT0 sheaf by an ample line bundle preserves the vanishing of higher cohomology (see [7, Thm. 2.3]).
@@ 7 Stability of the Ulrich Bundle
Because $V$ is stable on a curve, $End_{C}(V)=\mathbb{C}$ (see [12, Ch. 1]), hence $Hom_{A}(a_{*}V,a_{*}V)=\mathbb{C}$.
@@
Next we explain semihomogeneity. The sheaf $E=\Phi_{\mathcal{P}}(a_{*}V)$ is a Picard sheaf in the sense of Mukai (Fourier–Mukai transform of a sheaf pushed forward from a curve inside its Jacobian). Mukai proves that Picard sheaves on abelian varieties are semihomogeneous; see [3, §3–§4] (and also the exposition [14, Ch. 6]). Thus $E$ is semihomogeneous, and twisting by a line bundle preserves semihomogeneity, so $F=E(\Theta)$ is semihomogeneous as well.
Finally, by Mukai’s stability criterion, a simple semihomogeneous vector bundle on an abelian variety is $\mu_{H}$-stable with respect to every polarization $H$; see [5, Prop. 6.16] (or [14, Prop. 6.16]). Taking $H=\Theta$ yields that $F$ is $\mu_{\Theta}$-stable. ∎
@@ 8 Chern Class
with $\{\alpha_{i}\}$ a basis of $H^{1}(C)$ and $\{\beta_{i}\}$ the dual basis of $H^{1}(A)$ (see [14, Ch. 6]).
@@
A classical computation (see [14, Ch. 6] and [5, Prop. 6.16]) shows that
@@ 9.1 An upper bound for the Ulrich complexity
By Theorem 4.1, $F$ satisfies IT0. In particular, $F$ is continuously globally generated (CGG) in the sense of Pareschi–Popa (see [7, Remark 2.2(b)]).
@@ 10.2 A universal family after passing to a cover
This follows from the GIT construction of $M_{C}$ as a quotient of an open subset of a Quot scheme; see [12, Ch. 4]. Restricting to the stable locus and passing to a suitable cover yields a universal bundle on $C\times T$. ∎
@@ 10.3 Relative Fourier–Mukai transform
Openness of these properties follows from semicontinuity and cohomology and base change ([10, III.12], [7, §2]).
@@ 10.4 Relative evaluation and kernel construction
which is surjective on a dense open subset of $T$ by continuous global generation ([7, Remark 2.2]).
@@ 10.5 Stability in families
By Section 7, for general evaluation data the bundle $U(V_{t})$ is $\mu_{H}$-stable. Since stability is open in flat families ([12, Thm. 2.3.1]), there exists a dense open subset
@@ 10.7 Geometric consequences
Let $T^{\mathrm{st}}\subset T$ be the locus where $\mathcal{U}_{t}$ is $H$-stable. Since slope-stability is an open property in flat families of coherent sheaves (see [12, Thm. 2.3.1]), the subset $T^{\mathrm{st}}$ is Zariski open. By the stability results of Section 8 (for general evaluation data and sufficiently positive $m$), this locus is nonempty; hence $T^{\mathrm{st}}$ is dense in $T$.
@@
see [12, Cor. 4.6.6]. This proves the existence of $\pi$, $T^{\mathrm{st}}$, and $\Phi$ as stated.

### v2

# Ulrich Bundles on Jacobian Variety of a Curve

@@ 1 Introduction
Ulrich bundles occupy a central position in the interaction between algebraic geometry and commutative algebra. Originally introduced by Eisenbud and Schreyer in connection with linear determinantal representations and maximal Cohen–Macaulay modules [3]. Ulrich bundles are vector bundles charaterized by strong cohomological vanishing properties with respect to given polarization. they play a central role in study of syzygies and linear resolutions of projective varieties and Chow forms of projective varieties.
A fundamental conjecture of Eisenbud and Schreyer predicts that every polarized projective variety admits an Ulrich bundle. Despite substantial progress in special cases, the conjecture remains widely open. Ulrich bundles have been constructed for several classes of varieties, including many hypersurfaces and certain surfaces; see for example [1, 7]. However, the existence problem remains largely unexplored for irregular varieties and higher-dimensional abelian varieties.
In this paper we study the existence of Ulrich bundles on abelian varieties arising from Jacobians of smooth projective curves. Let $C$ be a smooth projective curve of genus $g\geq 2$, and let $A=J(C)$ denote its Jacobian equipped with the principal polarization $\Theta$. Our approach is based on Fourier–Mukai techniques and exploits the geometry of the Abel–Jacobi embedding $C\hookrightarrow J(C)$. Starting from a sufficiently positive line bundle $L$ on $C$, we consider the pushforward $a_{*}L$ along the Abel–Jacobi map and study its Fourier–Mukai transform. After a suitable twist by the principal polarization, we show that the resulting bundle on $J(C)$ satisfies strong regularity properties in the sense of Pareschi–Popa [8]. Using the theory of continuously globally generated bundles on abelian varieties, we construct a kernel bundle whose twist produces an Ulrich bundle on $(J(C),m\Theta)$ for sufficiently large $m$.
Our main result as follows.
@@
This construction verifies the Eisenbud–Schreyer conjecture for a large and geometrically natural class of abelian varieties, namely Jacobians of smooth projective curves endowed with sufficiently positive multiples of the principal polarization. To the best of our knowledge, this provides the first systematic construction of Ulrich bundles on Jacobians in arbitrary genus with respect to sufficiently positive polarization. It may be viewed as a higher-dimensional analogue of Beauville’s construction of Ulrich bundles on abelian surfaces [2].
Beyond the existence statement, the construction has further geometric consequences. The Ulrich bundles obtained here arise naturally from vector bundles associated to the curve and vary in positive-dimensional families. More precisely, the construction induces families of stable bundles on the Jacobian arising from moduli spaces of vector bundles on the curve.
@@
Fix integers $r\geq 2,d>r(2g-2)$, and let $M_{C}$ denote the moduli space of stable vector bundles on $C$ of rank $r$ and degree $d$. For $m\geq g+1$, let $M_{A}$ be the moduli space of $H$-stable bundles on $(A,m\Theta)$ with the numerical invariants of the Ulrich bundles constructed in Theorem 5.2 Then there exists a quasi-projective variety $T$, a generically finite surjective morphism
@@
These results illustrate a close relationship between the geometry of the curve and the moduli theory of vector bundles on its Jacobian. In particular, the construction connects Fourier–Mukai transforms, the geometry of Picard bundles, and the theory of Ulrich bundles. The techniques used here rely heavily on the modern theory of regularity on abelian varieties developed by Pareschi and Popa [8]. Their results on $IT_{0}$ sheaves and continuously globally generated bundles provide the key tools for controlling the cohomological behaviour of the Fourier–Mukai transforms appearing in our construction. We also make use of Mukai’s theory of semihomogeneous vector bundles [16].
Layout of the paper. In Section 2 we review some of the pre-requisites : Ulrich bundles, Abel-Jacobi map,Fourier–Mukai transforms, IT index properties, stable bundles and, semi-homogeneous bundles. We establish the $\mathrm{WIT}_{0}$ property and computes the rank of the transform in Section 3. In Section 4 we prove that $E(\Theta)$ satisfies $\mathrm{IT}_{0}$. Section 5 constructs Ulrich bundles and proves Theorem A. In Section 6 we discuss the numerical bound $m\geq g+1$. Section 7 studies slope stability. We compute Chern classes in Section 8. In Section 9 we specialize to line bundles and derive an upper bound for Ulrich complexity. We conclude Section 10 with the construction of the moduli morphism and proves the Theorem B.
Throughout the paper we work over the field $\mathbb{C}$.
@@ 2.2 The Abel–Jacobi embedding
A classical result (see [5, Ch. VII]) states that
@@ 2.7 Semihomogeneous bundles on abelian varieties
  • A semihomogeneous bundle is simple if and only if it is stable with respect to any polarization [16, Prop. 6.16].
  • If $\mathcal{E}$ is semihomogeneous, then its Chern classes are proportional to powers of $c_{1}(\mathcal{E})$ in the Néron–Severi group [16, Thm. 5.8].
@@ 5.1 Positivity and higher cohomology vanishing
For $k\geq 1$, the line bundle $\mathcal{O}_{A}(k\Theta)$ is ample. On an abelian variety, twisting an IT0 sheaf by an ample line bundle preserves the vanishing of higher cohomology (see [8, Thm. 2.3]).
@@ 5.3 Twisting and Ulrich vanishing
Infact, taking each $\alpha_{1},\alpha_{2},.....,\alpha_{N}$ non-trivial line bundles, we can conclude that
$H^{0}A,G(H-pH))=0\quad\forall p\geq 1$
Thus we conclude
$H^{i}(A,G(H-pH))=0\quad\forall i\geq 0,\forall p\geq 1$
@@ 7 Stability of the Ulrich Bundle
Because $V$ is stable on a curve, $End_{C}(V)=\mathbb{C}$ (see [13, Ch. 1]), hence $Hom_{A}(a_{*}V,a_{*}V)=\mathbb{C}$.
@@
Next we explain semihomogeneity. The sheaf $E=\Phi_{\mathcal{P}}(a_{*}V)$ is a Picard sheaf in the sense of Mukai (Fourier–Mukai transform of a sheaf pushed forward from a curve inside its Jacobian). Mukai proves that Picard sheaves on abelian varieties are semihomogeneous; see [4, §3–§4] (and also the exposition [15, Ch. 6]). Thus $E$ is semihomogeneous, and twisting by a line bundle preserves semihomogeneity, so $F=E(\Theta)$ is semihomogeneous as well.
Finally, by Mukai’s stability criterion, a simple semihomogeneous vector bundle on an abelian variety is $\mu_{H}$-stable with respect to every polarization $H$; see [6, Prop. 6.16] (or [15, Prop. 6.16]). Taking $H=\Theta$ yields that $F$ is $\mu_{\Theta}$-stable. ∎
@@ 8 Chern Class
with $\{\alpha_{i}\}$ a basis of $H^{1}(C)$ and $\{\beta_{i}\}$ the dual basis of $H^{1}(A)$ (see [15, Ch. 6]).
@@
A classical computation (see [15, Ch. 6] and [6, Prop. 6.16]) shows that
@@ 9.1 An upper bound for the Ulrich complexity
By Theorem 4.1, $F$ satisfies IT0. In particular, $F$ is continuously globally generated (CGG) in the sense of Pareschi–Popa (see [8, Remark 2.2(b)]).
@@ 10.2 A universal family after passing to a cover
This follows from the GIT construction of $M_{C}$ as a quotient of an open subset of a Quot scheme; see [13, Ch. 4]. Restricting to the stable locus and passing to a suitable cover yields a universal bundle on $C\times T$. ∎
@@ 10.3 Relative Fourier–Mukai transform
Openness of these properties follows from semicontinuity and cohomology and base change ([11, III.12], [8, §2]).
@@ 10.4 Relative evaluation and kernel construction
which is surjective on a dense open subset of $T$ by continuous global generation ([8, Remark 2.2]).
@@ 10.5 Stability in families
By Section 7, for general evaluation data the bundle $U(V_{t})$ is $\mu_{H}$-stable. Since stability is open in flat families ([13, Thm. 2.3.1]), there exists a dense open subset
@@ 10.7 Geometric consequences
Let $T^{\mathrm{st}}\subset T$ be the locus where $\mathcal{U}_{t}$ is $H$-stable. Since slope-stability is an open property in flat families of coherent sheaves (see [13, Thm. 2.3.1]), the subset $T^{\mathrm{st}}$ is Zariski open. By the stability results of Section 8 (for general evaluation data and sufficiently positive $m$), this locus is nonempty; hence $T^{\mathrm{st}}$ is dense in $T$.
@@
see [13, Cor. 4.6.6]. This proves the existence of $\pi$, $T^{\mathrm{st}}$, and $\Phi$ as stated.